OpenStudy (anonymous):

Another conversion from polar to rectangular form. r=4/(1+2sin(theta)) would that become 3y^2=x^2?

OpenStudy (sidsiddhartha):

16-16y those terms ??

OpenStudy (asnaseer):

@ddd231 - rather than just listing what you believe to be the answer it would be better if you listed your steps so we can help spot where you may have made a mistake

OpenStudy (dumbcow):

\[\sin \theta = \frac{y}{r}\] \[r (1 + \frac{2y}{r}) = 4\] \[r = 4 - 2y\] \[r^2 = 16 -16y +4y^2\] \[x^2 +y^2 = 16 -16y +4y^2\] \[x^2 = 3y^2 -16y +16\]

OpenStudy (anonymous):

i see what i did wrong now. when i multiplied over the 1+2sintheta i made the 4 into a 1

OpenStudy (anonymous):

Thanks for the answer @dumbcow and thanks d=for the help and you time as well @asnaseer @sidsiddhartha next time ill write out what i did

OpenStudy (asnaseer):

yw :)

OpenStudy (sidsiddhartha):

wlcm:)